OpenStudy (lilai3):

Write an equation in standard form of the line that passes through (2, -3) and (-3, 7). Thank you for your help! ***will reward medal /fan! ***

OpenStudy (lukecrayonz):

Do you know how to find slope

OpenStudy (lilai3):

yes; y2-y1 over x2-x1. correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now plug it in the formula y-y1=m(x-x1)

OpenStudy (lukecrayonz):

lilai, so find the slope

OpenStudy (lukecrayonz):

(7-(-3))/(-3-2)

OpenStudy (lilai3):

i still don't get it D:

OpenStudy (anonymous):

the equation of a line has the formula: y-y1=m(x-x1)

OpenStudy (lukecrayonz):

Okay lilai, find the slope and then tell me what it is.

OpenStudy (anonymous):

where m is the slope of the line

OpenStudy (anonymous):

m = change in y/ change in x = 7-(-3)/-3-2)=-2, so plugging in m you get, y-y1=-2(x-x1)

OpenStudy (anonymous):

choose any coordinate point from the question as your y1 and x1. so you get: y-(-3)=-2(x-2) y+3=-2(x-2)

OpenStudy (anonymous):

alternatively you can use the other coordinate point (-3,7) as your x1 and y1. So you will get y-7=-2(x-(-3) y-7=-2(x+3)

OpenStudy (lilai3):

hoohhhhhh thanks, you just clarified something for me! yay!

OpenStudy (anonymous):

y+3=-2(x-2) and y-7=-2(x+3) are both correct answer

OpenStudy (lilai3):

...if they are both the correct answer, then they are both standard form?

OpenStudy (anonymous):

yes you can choose whichever you want, both are correct

OpenStudy (anonymous):

but do simplify before giving the final answer

OpenStudy (lilai3):

oh! thank you. would you mind also answering a couple more? cuz 2mr i have a great big test :P if you don't mind of course. medal rewarded (;

OpenStudy (anonymous):

sure, but do simplify and put the answer in ax+by=c form

OpenStudy (anonymous):

since that is the standard form for a line

OpenStudy (lilai3):

Write an equation in slope-intercept form of the line satisfying the given conditions: is perpendicular to 4y=3x-8 and passes through (0,-6) THANKS AGAIN!

OpenStudy (anonymous):

Now first of all the slope intercept form of a line is given by the formula y=mx+c

OpenStudy (anonymous):

where m is the slope of the line and c is the y intercept

OpenStudy (lilai3):

sorry i started a new post so i could reward you another medal =) and also because they're too many comments on this one LOL i'll send you the link